# Bibliography.bib

### v1

@@ V-C The Feature Cloud Platform for Federated Learning
This platform takes advantage of federated learning and is an impressive work by Holzinger et al. [feature_cloud] presents a comprehensive exploration of the integration of domain knowledge graphs into deep learning for improved interpretability and explainability using Graph Neural Networks (GNNs). Federated Learning (FL) protects privacy by transmitting only model updates, reducing the risk of data breaches. Its decentralized approach enhances data security and compliance. FL promotes collaboration between devices, improving generalization with diverse data sets. In addition, it improves the learning algorithm by incorporating explanations and conceptual knowledge for better interpretability [li2020review]. The authors focus on using a protein-protein interaction (PPI) network to enrich deep neural networks for classification, enabling the detection of disease sub-networks using explainable AI. This work addresses the potential threats at each layer of the nested model and validates them accordingly.
@@
    Addressing Ethical Requirements: Privacy, data governance, and safety are addressed through federated learning approach. This use case accelerates biomedical research, which in turn benefits humanity and enables societal well-being. In addition, by decentralizing the training process, federated learning has the potential to be more environmentally friendly, making it a promising approach for reducing the carbon footprint of machine learning model training [qiu2023first].
@@
    Addressing the technical requirements: The framework included an expert-in-the-loop approach to develop AI workflows under the supervision of a domain expert, allowing for human agency and oversight. The domain expert evaluates data for biases related to non-discrimination and diversity. Knowledge graphs are implemented to incorporate prior domain knowledge. The potential threat at the data layer is validated. At the model layer, models are evaluated for fairness and technical robustness through continuous feedback from the domain expert. Open source models are preferred for training. Finally, the XAI is integrated to make the AI process transparent and trustworthy.
    In addition, this work also answers the extended XAI-QB in a summarized fashion. The questions answered are ”What was not done and why?”, ”What problems were solved?”, ”What was difficult?”, ”What did we learn?”, and ”What are the limitations and future results?
@@
GDPR, EU Guidelines for Trustworthy AI Expert-in-the-loop Domain expert assess data for bias related to non-discrimination and diversity.
Prior domain knowledge is incorporated in the form of knowledge graphs
Open source model for training, constant domain expert feedback to ensure fairness and robustness XAI techniques are implemented to enable transparency in decision-making
[…]
figureValidation of The Feature Cloud Platform through the Nested Model for AI design and validation
VI Discussion
We have discussed the potential threats and their validation methodologies, but we assume that the potential threats are a non-exhaustive list. There can be many potential threats and their validation methodologies that may exist at the given layer, and these potential threats and validation methodologies should be mapped to one of the layers of the nested model for AI design and validation.
Considering the need of the domain expert as an expert-in-the-loop, it is also possible that a single domain expert may not be able to list all the domain-specific requirements, model validation strategies, visualization techniques, etc., so we recommend multiple iterations of the nested model for AI design and validation. In addition, the XAI-QB should be kept in mind and the specific sub-type of the topic can be extended according to the requirements. For example, if a domain-specific data visualization is desired, then this question should be added to the XAI-QB in the data subtopic and then the appropriate methods should be used at the data level to address this question. The advantage of this framework is that it allows users to evaluate the AI workflow in human terms in collaboration with XAI-QB, which is an algorithmic question bank.
In addition, there may be questions such as ”Sometimes we don’t know the problem with the data until we start training”. One answer is that we are still at the data level because the issue with the data is still there in the training of the model. After the training of the model, it reflects that there is a problem with the data, then technically you are using the ML model to analyze the data for more potential threats, this will act as a methodology for identification of unknown or more potential threats. Accordingly, a question should be added to the XAI-QB under the data subtopic that asks, ”How do you identify unknown potential threats with the data that are not identifiable by normal bias identification methods?” The possibility of unstated biases is an open question, and some may be implicit threats.
The process of navigating the nested model for AI enables validation. We name this process as ”eliminating the potential threats of one layer through the functionality of the next layer”. Successfully navigating through each layer of the nested model layer for AI means that one has successfully resolved the threats present at each layer.
@@ VI Discussion
One could argue that questioning is not sufficient. However, formulating and asking the right questions brings up potential problems and assumptions from regulation all the way to predictions of an AI system. If a certain question or questions cannot be answered truthfully, or without reasonable doubt, the question(s) shall remain open to show transparency and potential pitfalls along every subsequent layer of the nested model for AI design and validation. Naturally certain questions require a hard limit, so to not proceeded to subsequent layers. This validation is meant to resolve any downstream problems when developing AI systems.
VII Conclusion
@@
In conclusion, our work is the first to bring together the disparate fields of AI and regulation. Our work establishes a common baseline for designing and validating AI under the umbrella of regulation, taking into account the complexity of the domain and the need for domain experts. Through this work, we are addressing discussions around the need to work at the intersection of AI and regulation to increase adoption in high-stakes domains such as healthcare. Adopting the nested model for AI design and validation will help realize the full potential of today’s state-of-the-art AI techniques in a complex, vast, but interconnected and globalized world.

### v2

@@ Introduction
This platform takes advantage of federated learning and is an impressive work by Holzinger et al. 24 presents a comprehensive exploration of the integration of domain knowledge graphs into deep learning for improved interpretability and explainability using Graph Neural Networks (GNNs). Federated Learning (FL) protects privacy by transmitting only model updates, reducing the risk of data breaches. Its decentralized approach enhances data security and compliance. FL promotes collaboration between devices, improving generalization with diverse data sets. In addition, it improves the learning algorithm by incorporating explanations and conceptual knowledge for better interpretability 27. The authors focus on using a protein-protein interaction (PPI) network to enrich deep neural networks for classification, enabling the detection of disease sub-networks using explainable AI. This work addresses the potential threats at each layer of the nested model and validates them accordingly (Refer figure: 10).
@@
    Addressing Ethical Requirements: Privacy, data governance, and safety are addressed through a federated learning approach. This use case accelerates biomedical research, which in turn benefits humanity and enables societal well-being. In addition, by decentralizing the training process, federated learning has the potential to be more environmentally friendly, making it a promising approach for reducing the carbon footprint of machine learning model training 39.
@@
    Addressing the technical requirements: The framework included an expert-in-the-loop approach to develop AI workflows under the supervision of a domain expert, allowing for human agency and oversight. The domain expert evaluates data for biases related to non-discrimination and diversity. Knowledge graphs are implemented to incorporate prior domain knowledge. The potential threat at the data layer is validated. At the model layer, models are evaluated for fairness and technical robustness through continuous feedback from the domain expert. Open-source models are preferred for training. Finally, the XAI is integrated to make the AI process transparent and trustworthy.
    In addition, this work also answers the extended XAI-QB in a summarized fashion. The questions answered are "What was not done and why?", "What problems were solved?", "What was difficult?", "What did we learn?", and "What are the limitations and future results?".
@@
Figure 10: Validation of The Feature Cloud Platform through the Nested Model for AI design and validation.
Discussion
We have discussed the potential threats and their validation methodologies, but we assume that the potential threats are a non-exhaustive list. There can be many potential threats and validation methodologies that may exist at the given layer, and these potential threats and validation methodologies should be mapped to one of the layers of the nested model for AI design and validation.
Considering the need for the domain expert as an expert-in-the-loop, it is also possible that a single domain expert may not be able to list all the domain-specific requirements, model validation strategies, visualization techniques, etc., so we recommend multiple iterations of the nested model for AI design and validation. In addition, the XAI-QB should be kept in mind and the specific sub-type of the topic can be extended according to the requirements. For example, if a domain-specific data visualization is desired, then this question should be added to the XAI-QB in the data subtopic and then the appropriate methods should be used at the data level to address this question. The advantage of this framework is that it allows users to evaluate the AI workflow in human terms in collaboration with XAI-QB, which is an algorithmic question bank.
In addition, there may be questions such as "Sometimes we don’t know the problem with the data until we start training". One answer is that we are still at the data level because the issue with the data is still there in the training of the model. After the training of the model, it reflects that there is a problem with the data, then technically you are using the ML model to analyze the data for more potential threats, this will act as a methodology for the identification of unknown or more potential threats. Accordingly, a question should be added to the XAI-QB under the data subtopic that asks, "How do you identify unknown potential threats with the data that are not identifiable by normal bias identification methods?" The possibility of unstated biases is an open question, and some may be implicit threats.
The process of navigating the nested model for AI enables validation. We name this process as "eliminating the potential threats of one layer through the functionality of the next layer". Successfully navigating through each layer of the nested model layer for AI means that one has successfully resolved the threats present at each layer.
@@ Discussion
The nested model takes into account both the AI in production and the AI in development. AI regulations are designed to address both the development and production or deployment phases of AI systems. They aim to create a framework that promotes innovation while safeguarding against potential risks and ensuring ethical use of AI technology throughout its lifecycle. One of the examples that aligns with this scenario is the EU AI Act. It takes a comprehensive approach that covers the entire AI lifecycle. The legislation establishes rules for AI developers, deployers, and end-users. It also sets up governance structures at the national and international levels to oversee AI regulation. Furthermore, it creates mechanisms for ongoing assessment and adaptation of regulations as AI technology evolves. For instance, for AI in development, transparency is required in the AI model’s design and training process. Additionally, the results and decision-making processes must be communicated effectively with the help of audience-centric XAI techniques.
There exist various regulations for the domains. One such example is pubmed regulations for the biomedical domain. Two seminal works focus on guidelines and best practices for health informatics and software development in healthcare, with a particular emphasis on artificial intelligence (AI) applications 20; 21. These articles highlight the importance of creating reproducible and reusable biomedical software, fostering technology transfer in health informatics, and leveraging AI’s potential in medical research and precision medicine. Furthermore, the articles address the software life cycle in health informatics and the role of AI in supporting clinical decision-making, which can enhance patient monitoring, diagnostics, and prognostics. They also discuss crucial aspects often considered in healthcare software regulations, such as ensuring software reliability, maintaining data privacy and security, validating AI algorithms for clinical use, and implementing proper documentation and version control. It is crucial to acknowledge that the specific regulations pertaining to health informatics and AI in healthcare may vary by country and jurisdiction. Therefore, it is essential for developers and healthcare providers to consult with the relevant regulatory bodies to ensure that they are aware of the most up-to-date and applicable regulations when implementing health informatics solutions.
@@
One could argue that questioning is not sufficient. However, formulating and asking the right questions brings up potential problems and assumptions from regulation all the way to predictions of an AI system. If a certain question or questions cannot be answered truthfully, or without reasonable doubt, the question(s) shall remain open to show transparency and potential pitfalls along every subsequent layer of the nested model for AI design and validation. Naturally, certain questions require a hard limit, so do not proceed to subsequent layers. This validation is meant to resolve any downstream problems when developing AI systems.
The future work would focus on developing an audience-centric XAI that takes into account the background of the end-users. The development of audience-centric XAI necessitates a comprehensive approach that prioritizes user needs and contexts. Key requirements include the implementation of user-centric design principles to tailor explanations for diverse user groups, the provision of real-time and actionable insights, and the utilization of scenario-based methods to elicit specific user requirements. The integration of human-computer interaction principles is of paramount importance for the creation of effective explanation interfaces. Furthermore, the assurance of compliance with legal and regulatory standards is of critical significance. By meeting these requirements, XAI systems can deliver customized, meaningful explanations that enhance user understanding and trust across various contexts, while also fulfilling legal obligations. This approach aims to bridge the gap between complex AI systems and end-users, making AI decisions more transparent and interpretable for different audiences. 30 It is not uncommon for there to be debate surrounding the provision of overly detailed explanations, which could potentially be exploited or manipulated by malicious actors. For instance, counterfactual explanations may be employed to identify adversarial samples with greater ease. Despite this, XAI systems remain susceptible to the biases inherent in the data and algorithms utilized, given that humans set the parameters for these systems 41. One potential solution is to involve multiple domain experts in lieu of a single expert.
Limitations of study
The achievement of international harmonization in AI regulation represents a significant challenge, given the diverse regulatory approaches that exist across countries and regions. Nevertheless, there are indications of progress, including a growing consensus on high-risk applications and various international initiatives working towards harmonization. The key challenges that must be addressed include the definition of AI, the balancing of innovation with risk mitigation, and the navigation of international tensions. On a more positive note, there is an increasing political will for regulation, signs of trans-Atlantic convergence, and the potential influence of EU regulations globally. Potential solutions include the establishment of global governance institutions, the adaptation of best practices from other high-tech sectors, and the development of open-source AI software. While full harmonization remains a challenging objective, there is a growing consensus towards more coordinated approaches to AI regulation through continued dialogue, cooperation, and flexible frameworks adaptable to this rapidly evolving technology. Our nested model is designed to facilitate international harmonization for AI regulations. The essential requirements identified by regulatory bodies in different countries may vary, but the underlying fundamental principle to satisfy these requirements remains consistent.
Conclusion
@@
In conclusion, our work is the first to bring together the disparate fields of AI and regulation. Our work establishes a common baseline for designing and validating AI under the umbrella of regulation, taking into account the complexity of the domain and the need for domain experts. Through this work, we are addressing discussions around the need to work at the intersection of AI and regulation to increase adoption in high-stakes domains such as healthcare. Adopting the nested model for AI design and validation will help realize the full potential of today’s state-of-the-art AI techniques in a complex, vast, but interconnected, and globalized world.
Author Contributions
Conceptualization, A.D., G.H.; Methodology. A.D., G.H.; Writing – Original Draft, A.D., Z.Y., and G.H.; Writing – Review & Editing, A.D., Z.Y., and G.H.; Supervision, Z.Y. and G.H., Project Administration, G.H.
Declaration of Interests
The authors declare no competing interests.
@@
[…]
Resource availability
Lead contact
Further information and requests should be directed to and will be fulfilled by the lead contact, Akshat Dubey (DubeyA@rki.de)
Materials availability
Given the nature of the research, no datasets or codes are used or produced.
Data and code availability
No data and code are involved in the research conducted.
Method details
The research is focused on the evolving landscape of AI regulations. Although artificial intelligence (AI) has experienced significant growth in recent years, it has yet to achieve its full potential in real-world applications. This research is divided into four sections. The initial section examines the advantages of AI and elucidates the reasons behind the relatively slow rate of adoption of AI in critical and high-stakes domains. The advent of AI has brought with it a number of significant ethical and legal concerns. Consequently, there is a pressing need to address not only the regulatory policies that will facilitate the implementation of AI in real-world use cases but also how practitioners design and validate AI applications and workflows. The second part of the research is concerned with the potential of an audience-centric approach to AI, in conjunction with explainable artificial intelligence, to address significant concerns with AI regulations and satisfy the guidelines set out by regulatory authorities. The third section of the research presents the XAI-QB, which is a question bank comprising algorithm-informed questions. These questions serve to assess and validate the AI, thereby fostering trust, faithfulness, and transparency in AI through the application of design principles. In the fourth part, we have developed a nested model for AI design and validation on top of XAI-QB. The nested model contains five distinct layers: regulations, domain, data, model, and predictions. At each level, specific prototypical questions must be addressed, with due consideration given to the guidelines set forth by regulatory authorities. Additionally, illustrative examples have been provided.